# references.bib

### v1

@@ Abstract
Compositional generalization (the ability to respond correctly to novel combinations of familiar components) is thought to be a cornerstone of intelligent behavior. Compositionally structured (e.g. disentangled) representations are essential for this; however, the conditions under which they yield compositional generalization remain unclear. To address this gap, we present a general theory of compositional generalization in kernel models with fixed, potentially nonlinear representations (which also applies to neural networks in the “lazy regime”). We prove that these models are functionally limited to adding up values assigned to conjunctions/combinations of components that have been seen during training (“conjunction-wise additivity”), and identify novel compositionality failure modes that arise from the data and model structure, even for disentangled inputs. For models in the representation learning (or “rich”) regime, we show that networks can generalize on an important non-additive task (associative inference), and give a mechanistic explanation for why. Finally, we validate our theory empirically, showing that it captures the behavior of deep neural networks trained on a set of compositional tasks. In sum, our theory characterizes the principles giving rise to compositional generalization in kernel models and shows how representation learning can overcome their limitations. We further provide a formally grounded, novel generalization class for compositional tasks that highlights fundamental differences in the required learning mechanisms (conjunction-wise additivity).
@@ 1 Introduction
Humans’ understanding of the world is inherently compositional: once familiar with the concepts “pink” and “elephant,” we can immediately imagine a pink elephant. Stitching together concepts in this way allows humans to generalize far beyond our prior experience, preparing us to cope with unfamiliar situations and imagine things that do not yet exist [lake_building_2017, frankland_concepts_2020]. Understanding the basis of compositional generalization in humans and animals, and building it into machine learning models, is a long-standing and historically vexing problem [fodor_connectionism_1988, battaglia_relational_2018, lake_generalization_2018, hupkes_compositionality_2020, keysers_measuring_2020]. Accordingly, a broad range of studies have investigated the conditions under which compositionally structured (i.e. “disentangled”) representations can be learned [hinton_transforming_2011, bengio_representation_2013, higgins_beta-vae_2017, locatello_challenging_2019, esmaeili_structured_2019, trauble_disentangled_2021, whittington_disentanglement_2023]. However, it remains unclear whether learning these representations is actually useful: while some work suggests that disentangled representations improve compositional generalization [esmaeili_structured_2019, tokmakov_learning_2019, van_steenkiste_are_2019, whittington_constellation_2021], other studies challenge this view [locatello_challenging_2019, montero_role_2021, schott_visual_2022, xu_compositional_2022]. In particular, it is often unclear how insights on a certain compositional task generalize to others, as the relationship among them remains unclear [hupkes_compositionality_2020].
To clarify when compositionally structured representations yield compositional generalization, we define a family of compositional tasks, theoretically analyse the compositional generalization of kernel models, and then validate our theory in several relevant deep neural network architectures. Kernel models are an important class of statistical models in their own right (describing e.g. linear readout fine-tuning), provide a simplified approximation to neural network learning under certain conditions [jacot_neural_2018, chizat_lazy_2019], and more broadly have provided fundamental insight into generalization behavior in humans and machine learning models [jakel_generalization_2008, jakel_does_2009, canatar_spectral_2021, canatar_out--distribution_2021].
@@
Despite the broad importance and relative simplicity of kernel models, it remains largely unclear how they generalize on compositional tasks [[, though see ]; see Section 2]abbe_generalization_2023,lippl_mathematical_2024. To address this gap, we present a theory of compositional generalization in kernel models. We define a new generalization class for compositional tasks and show that characterizing kernel models on such tasks helps us understand realistic neural network architectures. Our specific contributions are as follows:
@@
  • We prove fundamental limits on the generalization behavior of kernel models with disentangled inputs, showing that they are constrained to summing up values implicitly assigned to each component or combination of components seen during training. We define the class of compositional tasks with this form as “conjunction-wise additive” (Section 4.1).
  • We define a metric to quantify how salient different component “conjunctions” (entangled features unique to component combinations) are in the representation (Section 4.2).
@@
  • For conjunction-wise additive tasks, we then highlight two important failure modes impacting generalization (memorization leak and shortcut bias) (Section 4.3), which describe how biases in the training data can limit compositional generalization even for kernel models with disentangled inputs.
@@
  • We validate our theory in several deep neural network architectures, showing that it captures their behavior on conjunction-wise additive tasks (Section 4.4).
  • Finally, we show how feature learning can overcome certain limitations of kernel models and learn non-additive tasks (Section 4.5).
@@ 2 Related work
Compositional generalization. Compositionality is an important theme across human and machine reasoning problems, including visual reasoning [lake_human-level_2015, johnson_clevr_2017, marino_ok-vqa_2019, schwartenbeck_generative_2023, zhou_compositional_2024], language production [hupkes_compositionality_2020], rule learning [ito_compositional_2022, abdool_continual_2023], image generation [okawa_compositional_2023], and robotics [zhou_policy_2022]. Recent breakthroughs (especially in language and computer vision) have led to massive improvements in models’ compositional capacities, but in some cases, these models still fail spectacularly [srivastava_beyond_2023, lewis_does_2023, west_advances_2023, ma_crepe_2023, chen_skills--context_2023]. Attempts to improve compositional generalization often leverage meta-learning [lake_human_2019, mitchell_challenges_2021, kumar_meta-learning_2021, wu_adaptive_2023, lake_human-like_2023], or modular architectures, in the hopes that different modules will specialize for different components [andreas_neural_2017]. End-to-end training of these architectures often does not result in the desired modular specialization [bahdanau_systematic_2019, mittal_is_2022, jarvis_specialization_2023]; notably, [schug_discovering_2024] theoretically clarify the necessary conditions for correct modular specialization and compositional generalization in hypernetworks.
Here we highlight that kernel models add up values assigned to each combination of features (or “conjunction”) seen during training. Notably, constraining a network to be additive can be sufficient for compositional generalization [lachapelle_additive_2023, wiedemer_compositional_2023]. Even without such constraints, language models encode many semantic concepts in an additive manner [mikolov_distributed_2013, naito_revisiting_2021]. Conjunctive codes (representations specific to a particular feature combination), on the other hand, have a long history in neuroscience [alvarado_properties_1992, baker_impact_2002], and are theoretically linked to forming highly specific, episodic memories.
Kernel and rich regime. Prior work has revealed two key strategies by which deep neural networks learn [chizat_lazy_2019, woodworth_kernel_2020]. In the kernel (or “lazy”) regime (which can be induced with large initial weights or wide networks), the networks’ learning is well approximated by gradient descent on a model with a fixed representation [jacot_neural_2018]. In the rich regime (brought forth, for example, by small initial weights, small width, low-rank connectivity, or loss functions like cross-entropy), the networks learn structured (i.e. abstract and sparse) representations over the course of learning [savarese_how_2019, woodworth_kernel_2020, chizat_implicit_2020, lyu_gradient_2020, saxe_neural_2022, liu_how_2023, domine_exact_2023]. This is often suggested to give rise to better generalization and more human-like behavior in neural networks [fort_deep_2020, vyas_limitations_2022, flesch_orthogonal_2022].
@@
When using gradient descent, ridge regression, or similar learning algorithms to train the readout weights of a model $f_{w}(x)=w^{T}r(x)$, this model depends on its representation $r(x)$ only through its induced kernel $K(x,x^{\prime})=\langle r(x),r(x^{\prime})\rangle$ [scholkopf_kernel_2000]. Specifically, when trained on a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, it can be written in its “dual form” as
$f_{a}(x)=\sum_{i=1}^{n}a_{i}K(x,x_{i}).$ (1)
@@
Gradient descent and ridge regression learn the readout weights that describe the training data with minimal $\ell_{2}$-norm; this is also true of neural networks in the kernel regime [soudry_implicit_2018, gunasekar_characterizing_2018, ji_gradient_2020]. Norm minimization is a standard theoretical framework for analyzing how representational geometry influences generalization [canatar_spectral_2021, canatar_spectral_2023] and we here apply it to the compositional task setting. This is similar to [lippl_mathematical_2024] who characterize model behavior on a specific compositional task (transitive ordering) and [abbe_generalization_2023] who characterize the inductive bias of norm minimization for inputs with binary components in the limit of infinite components. Compared to [abbe_generalization_2023, lippl_mathematical_2024], we analyze a broader range of compositional tasks and derive exact constraints for finite numbers of components.
[…]
Shortcut learning refers to models exploiting spurious correlations between certain features and the target to learn the training data and has been essential for understanding why the performance of deep neural networks can drop substantially when tested out-of-distribution [geirhos_shortcut_2020]. Models are more likely to use shortcut features that are easily extracted from the input and highly predictive of the output [hermann_what_2020, tartaglini_developmentally-inspired_2022, hermann_foundations_2023]. Shortcut learning has also been connected to the implicit bias of gradient descent [shah_pitfalls_2020, nagarajan_understanding_2021]. Our findings highlight a particular compositional shortcut: models tend to use a more salient, partially predictive component conjunction (memorizing the remaining data) over the fully predictive but less salient ground-truth conjunction.
@@ 3.1 Task space
We assume that our model extracts a set of components $z=\sum_{c=1}^{C}z_{c}$ from its input $x$. Each $z_{c}\in Z_{c}\subset\mathbb{R}^{d}$ is drawn from a discrete set of possible components. For example, $z$ could specify the color and shape of an object, and the background color of an image (Fig. 1a). Alternately, $z$ may express different words in an input sentence (“the red sphere is in front of the blue background”). We assume that each component $z_{c}$ is represented in an orthogonal subspace (e.g. by concatenating representations of individual components, Fig. 1a). Further, because we consider categorical components, we can assume that for each component $c$, all representations $z_{c}\in Z_{c}$ have equal magnitude $m_{c}$ and are represented with equal similarity $\rho_{c}$ to alternative component values. The target, $y\in\mathbb{R}$, is some function of $z$ (see Section 3.3). After training models on certain component combinations $Z^{\text{train}}\subset Z=\prod_{c=1}^{C}Z_{c}$, we assess generalization on all other combinations, $Z^{\text{test}}:=Z\setminus Z^{\text{train}}$ to understand how biases in the training data affect compositional generalization.

### v2

@@ Abstract
Compositional generalization (the ability to respond correctly to novel combinations of familiar components) is thought to be a cornerstone of intelligent behavior. Compositionally structured (e.g. disentangled) representations are essential for this; however, the conditions under which they yield compositional generalization remain unclear. To address this gap, we present a general theory of compositional generalization in kernel models with fixed representations, a tractable framework for characterizing the impact of dataset statistics on generalization. We find that kernel models are constrained to adding up values assigned to each combination of components seen during training (“conjunction-wise additivity”). This imposes fundamental restrictions on the set of tasks these models can learn, in particular preventing them from transitively generalizing equivalence relations. Even for compositional tasks that kernel models can in principle learn, we identify novel failure modes in compositional generalization that arise from biases in the training data and affect important compositional building blocks such as symbolic addition and context dependence (memorization leak and shortcut bias). Finally, we empirically validate our theory, showing that it captures the behavior of deep neural networks (convolutional networks, residual networks, and Vision Transformers) trained on a set of compositional tasks with similarly structured data. Ultimately, this work provides a theoretical perspective on how statistical structure in the training data can affect compositional generalization, with implications for how to identify and remedy failure modes in deep learning models.
@@ 1 Introduction
Humans’ understanding of the world is inherently compositional: once familiar with the concepts “pink” and “elephant,” we can immediately imagine a pink elephant. Stitching together concepts in this way lets humans generalize far beyond our prior experience, allowing us to cope with unfamiliar situations and imagine things that do not yet exist (Lake et al., 2017; Frankland & Greene, 2020). Understanding the basis of compositional generalization in humans and animals, and building it into machine learning models, is a long-standing and historically vexing problem (Fodor & Pylyshyn, 1988; Battaglia et al., 2018; Lake & Baroni, 2018). While a wide range of studies have investigated the conditions under which compositionally structured (i.e. “disentangled”) representations can be learned (Hinton et al., 2011; Higgins et al., 2017; Träuble et al., 2021; Whittington et al., 2023; Ren et al., 2023), it remains unclear when learning these representations is actually useful to a downstream neural network. Some work suggests that disentangled representations improve compositional generalization (Esmaeili et al., 2019; van Steenkiste et al., 2019). However, others challenge this view (Locatello et al., 2019; Schott et al., 2022). In general, it remains unclear why insights from some compositional tasks fail to generalize to others, and a systematic understanding of the relationships among compositional tasks remains elusive (Hupkes et al., 2020).
To make progress on this question, we focus on standard statistical learning, a fundamental basis for compositional generalization that affects almost any machine learning model. Generalization in statistical learning depends on the similarities between different inputs. Notably, in compositionally structured representations, inputs that have components in common (say, a red circle and a blue circle) are more similar to each other than inputs that do not (say, a red circle and a blue square). As a result, the compositional structure of a task is reflected in its dataset statistics, influencing the generalization behavior of the model.
To understand when and how statistical learning leads to successful compositional behavior, we developed a theory for the behavior of kernel models on compositional tasks. Kernel models are an important class of statistical models that are able to provide a simplified approximation to neural networks while maintaining analytical tractability (Jäkel et al., 2008, 2009). For example, kernel models accurately describe the behavior of learning and fine-tuning in neural networks under certain conditions (Jacot et al., 2018; Malladi et al., 2023). More broadly, they provide a tractable framework for characterizing the impact of dataset statistics on generalization (Canatar et al., 2021a, b).
@@
Despite their broad relevance and relative simplicity, it has remained unclear how even kernel models generalize on compositional tasks (though see Abbe et al., 2023; Lippl et al., 2024, see Section 2). To address this gap, we present a theory of compositional generalization in kernel models. We define a general family of compositional tasks with no constraints on the input-output mapping (ensuring broad applicability of our theory), theoretically characterize the compositional behavior of kernel models trained on such tasks, and go beyond kernel models to empirically validate our theory in several relevant deep neural network architectures. Our specific contributions are as follows:
@@
  • In Section 4, we show that kernel models with disentangled inputs are constrained to summing up values implicitly assigned to each component or combination of components seen during training. We call such computations “conjunction-wise additive.”
@@
  • In Section 5, we then characterize how representational geometry determines whether kernel models will generalize successfully on conjunction-wise additive compositional tasks, highlighting two important failure modes (memorization leak and shortcut bias). This highlights that disentangled representations are not sufficient for downstream compositional generalization (even on conjunction-wise additive tasks) and explains why.
@@
  • Finally, in Section 6, we validate our theory in several deep neural network architectures, showing that it captures their behavior on conjunction-wise additive tasks.
Overall, we take a step towards a general theory of compositional generalization by characterizing the impact of dataset statistics on compositional generalization for a broad range of tasks. Our theory systematically clarifies the compositional generalization behavior of deep neural networks and lays a foundation for the design of new learning mechanisms that overcome their limitations.
@@ 2 Related work
Compositional generalization. Compositionality is an important theme in both human and machine reasoning problems, including visual reasoning (Lake et al., 2015; Johnson et al., 2017; Schwartenbeck et al., 2023), language production (Hupkes et al., 2020), and rule learning (Ito et al., 2022; Abdool et al., 2023). Recent breakthroughs have led to massive improvements in models’ compositional capacities, but in some cases, these models still fail spectacularly (Srivastava et al., 2023; Lewis et al., 2023; West, 2023). Attempts to improve compositional generalization often leverage meta-learning (Mitchell et al., 2021; Wu et al., 2023; Lake & Baroni, 2023) or modular architectures (Andreas et al., 2017), in the hopes that different modules will specialize for different components (Schug et al., 2024). Constraining a network to be linear (or additive) can guarantee modular specialization and compositional generalization (Lachapelle et al., 2023; Wiedemer et al., 2023a, b). However, these networks are extremely limited in the tasks they can learn (as we will show below) and end-to-end training of modular architectures without such constraints often does not result in the desired specialization (Bahdanau et al., 2019; Mittal et al., 2022; Jarvis et al., 2023).
Conjunctions. Henceforth, we will call a combination of components a conjunction. Conjunctive codes have a long history in neuroscience (Alvarado & Rudy, 1992; Baker et al., 2002), and are theoretically linked to the formation of highly specific episodic memories.
@@
Kernel regime. When using gradient descent, ridge regression, or similar learning algorithms to train the readout weights of a model $f_{w}(z)=w^{T}\phi(z)$, it can be shown that the resulting behavior of that model can be described in terms of the kernel $K(z,z^{\prime})=\phi(z)^{T}\phi(z^{\prime})$ induced by the representation $\phi(z)$ (Schölkopf, 2000). We will refer to such models as kernel models. Notably, when trained on a dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, a kernel model can be written in its “dual form” as
$f_{a}(z)=\sum_{i=1}^{n}a_{i}K(z,z_{i}),\mbox{ for dual coefficients}a\in\mathbb{R}^{n}.$ (1)
Prior work has shown that neural networks with large initial weights or wide architectures are well approximated by gradient descent on a model with a fixed representation (a kernel model) (Jacot et al., 2018). This is called the “kernel regime,” in contrast to the feature-learning regime, which is brought forth, for example, by small initial weights or small width (Chizat et al., 2019).
@@
Norm minimization. Gradient descent, ridge regression, and neural networks in the kernel regime learn the readout weights that describe the training data with minimal $\ell_{2}$-norm (Soudry et al., 2018; Gunasekar et al., 2018; Ji et al., 2020). Norm minimization is a standard theoretical framework for analyzing how representational geometry and dataset statistics influence generalization (Canatar et al., 2021b, 2023) and we here apply it to the compositional task setting. This is similar to Lippl et al. (2024) who characterize model behavior on a specific compositional task (transitive ordering) and Abbe et al. (2023) who characterize the inductive bias of norm minimization for inputs with binary components in the limit of infinite components. Compared to these prior works, we analyze a broader range of compositional tasks and derive exact constraints for finite numbers of components.
Memorization leak. Machine learning models sometimes memorize (parts of) their training data instead of learning a generalizable rule (Zhang et al., 2020; Dasgupta et al., 2022). This may improve generalization on long-tailed data (Feldman, 2020), but often impairs out-of-distribution performance (Elangovan et al., 2021). We find that models partially memorize their training data even when they extract the correct rule. Jarvis et al. (2023) analyze a related phenomenon in deep linear networks.
Shortcut learning. Shortcut learning refers to models exploiting spurious correlations between certain features of the input data and the target (Shah et al., 2020; Nagarajan et al., 2021). This substantially impacts their performance out of distribution (where those correlations may not hold) (Geirhos et al., 2020). We theoretically analyze how shortcut biases impact compositional generalization.
@@ 3.1 Task space
We assume that our model extracts from its input a set of disentangled components $(z_{c})_{c=1}^{C}$, e.g. color, shape, and background (Fig. 1a). For instance, the model could have been pre-trained to extract these components from an image or an input sentence. Each component $z_{c}\in Z_{c}\subset\mathbb{R}^{d}$ is drawn from a discrete set of possible components represented by $d$-dimensional real vectors and because we consider categorical components, we assume that all components have unit magnitude and are orthogonal to each other (these assumptions can be relaxed, Section A.1). We then assume that the model represents these components by adding them together: $z:=\sum_{c=1}^{C}z_{c}$. A simple example of this is given by a concatenation of one-hot representations of all components (Fig. 1a). The target, $y\in\mathbb{R}$, is given by an arbitrary function of $z$, ensuring that our framework is agnostic to the underlying compositional structure. After training models on certain combinations of components $Z^{\text{train}}\subset Z=\prod_{c=1}^{C}Z_{c}$, we assess generalization on all other combinations $Z^{\text{test}}:=Z\setminus Z^{\text{train}}$.
Our theory applies to any task in this family; that is, any task with a fixed number of categorical input components and a scalar output, regardless of the input-output mapping. Below we describe three example tasks within this general framework that represent important building blocks for compositional reasoning in machine learning and cognitive science. We will focus on these tasks in the main text, to illustrate our general theory and analyze its consequences in more detail. In Appendix D, we describe additional tasks captured by our theory.
Symbolic addition. Many tasks involve inferring a magnitude associated with different underlying components (e.g. adding handwritten digits) (Lorenzi et al., 2021; Sheahan et al., 2021). We consider two components $(z_{1},z_{2})$ with unobserved assigned values $v_{1}(z_{1})$ and $v_{2}(z_{2})$. The target is the sum of those values: $y=v_{1}(z_{1})+v_{2}(z_{2})$. After sufficient exposition to individual items, a model with an additive structure can generalize to novel combinations of items. In particular, we consider nine input elements $[-4],[-3],\dotsc,[4]$ with associated values $-4,-3,\dotsc,4$ and training sets containing all pairs where at least one component is equal to a certain subset of values (given by the rows and columns in the training set in Fig. 1b). We vary the size of this subset (i.e. the number of rows/columns) as well as whether the task requires only interpolation or also extrapolation.
Context dependence. The relevance of different stimuli often depends on our current context. Taking this into account is a crucial aspect of cognition in humans and animals (Bouton et al., 1999; Taylor & Ivry, 2013; Parker & Hollister, 2014; Ito et al., 2022). We therefore consider a task with three input components $(z_{\text{co}},z_{f1},z_{f2})$. The context ($z_{\text{co}}$) has two possible values specifying whether $z_{f1}$ or $z_{f2}$ determine the response. Both features have six possible values which are split up in two categories. If the model has learned this context dependence, it should be able to generalize to novel feature combinations. We evaluate on the subset of data for which $z_{f1}$ indicates Cat. 2 and $z_{f2}$ indicates Cat. 1. In the most extreme generalization test (CD-3), we leave out the entire subset; in easier versions, we leave out combinations of two or one of those features (CD-2 and CD-1) (Fig. 1c).
Transitive equivalence. Relational reasoning is an important example of compositional generalization and often involves extending learned relations to new item combinations (Halford et al., 2010; Battaglia et al., 2018). Given an unobserved (and arbitrary) equivalence relation, the task is to determine whether two presented items $(z_{1},z_{2})$ are equivalent. The model should generalize to novel item pairs using transitivity ($A=B$ and $B=C$ imply $A=C$) (Fig. 1d). This is an important instance of relational cognition (often studied as “associative inference” in cognitive science (Schlichting & Preston, 2015; Spalding et al., 2018)). Although prior work has found that kernel models often successfully generalize on a transitive ordered relation ($A>B$ and $B>C$ imply $A>C$, Lippl et al., 2024), the behavior of kernel models on equivalence relations has remained unclear.